# Large Language Model-Brained GUI Agents: A Survey

### v9

@@ 8 Models for Optimizing LLM-Brained GUI Agents
TABLE XVIII: Overview of foundation models for LLM-brained GUI agents.
@@ 8.1 Foundation Models
Foundation models serve as the core of LLM-powered GUI agents, providing the essential capabilities for understanding and interacting with graphical user interfaces. Recent advancements in both close-source and open-source MLLMs have significantly enhanced the potential of GUI agents, offering improvements in efficiency, scalability, and multimodal reasoning. This subsection explores these foundation models, highlighting their innovations, contributions, and suitability for GUI agent applications. For a quick reference, Table XVIII presents an overview of the key models and their characteristics.
@@ 8.2 Large Action Models
TABLE XIX: An overview of GUI-optimized models on web platforms.
@@ 8.3 LAMs for Web GUI Agents
In the domain of web-based GUI agents, researchers have developed specialized LAMs that enhance interaction and navigation within web environments. These models are tailored to understand the complexities of web GUIs, including dynamic content and diverse interaction patterns. We present an analysis of LAMs tailored for web GUI agents in Table XIX.
@@
TABLE $X$: An overview of GUI-optimized models on mobile platforms.
@@ 8.4 LAMs for Mobile GUI Agents
Mobile platforms present unique challenges for GUI agents, including diverse screen sizes, touch interactions, and resource constraints. Researchers have developed specialized LAMs to address these challenges, enhancing interaction and navigation within mobile environments. We present an overview of LAMs tailored for mobile GUI agents in Table XX.
@@
These models collectively advance the field of mobile GUI agents by addressing platform-specific challenges through innovative training methods, efficient model architectures, and specialized datasets.
TABLE XXI: An overview of GUI-optimized models on computer platforms.
@@ 8.5 LAMs for Computer GUI Agents
For desktop and laptop environments, GUI agents must handle complex applications, multitasking, and varied interaction modalities. Specialized LAMs for computer GUI agents enhance capabilities in these settings, enabling more sophisticated task execution. We overview of LAMs for computer GUI agents across in Table XXI.
@@
TABLE XXII: An overview of GUI-optimized models on cross-platform agents.
@@ 8.6 Cross-Platform Large Action Models
To achieve versatility across various platforms, cross-platform LAMs have been developed, enabling GUI agents to operate seamlessly in multiple environments such as mobile devices, desktops, and web interfaces. We provide an analysis of LAMs tailored for cross-platform GUI agents in Table XXII.
@@ 9 Evaluation for LLM-Brained GUI Agents
In the domain of GUI agents, evaluation is crucial for enhancing both functionality and user experience [54, 56] and should be conducted across multiple aspects. By systematically assessing these agents’ effectiveness across various tasks, evaluation not only gauges their performance in different dimensions but also provides a framework for their continuous improvement [447]. Furthermore, it encourages innovation by identifying areas for potential development, ensuring that GUI agents evolve in tandem with advancements in LLMs and align with user expectations.
@@ 9.1 Evaluation Metrics
Evaluating GUI agents requires robust and multidimensional metrics to assess their performance across various dimensions, including accuracy, efficiency, and compliance (e.g., safety). In a typical benchmarking setup, the GUI agent is provided with a natural language instruction as input and is expected to autonomously execute actions until the task is completed. During this process, various assets can be collected, such as the sequence of actions taken by the agent, step-wise observations (e.g., DOM or HTML structures), screenshots, runtime logs, final states, and execution time. These assets enable evaluators to determine whether the task has been completed successfully and to analyze the agent’s performance. In this section, we summarize the key evaluation metrics commonly used for benchmarking GUI agents. Note that different research works may use different names for these metrics, but with similar calculations. We align their names in this section.
@@
    Task Success Rate: Task success rate measures the successful task completion over all tasks set in the benchmark. It evaluates whether the final task completion state is achieved while ignoring the intermediate steps. This metric provides an overall measure of end-to-end task completion, reflecting the agent’s ability to handle complex workflows holistically [366, 370, 357].
@@ 9.2 Evaluation Measurements
    • Partial or Fuzzy Match: Allows for approximate matches, which are useful for handling minor variations such as typos or synonyms.
@@
    • System Controls and APIs: Ensuring correct interaction with operating system controls or APIs, such as file dialogs, system menus, or notifications in desktop environments.
@@
Each of these measurement techniques contributes to a comprehensive evaluation framework, ensuring that the agent not only completes tasks but does so with precision, efficiency, and adaptability. Together, they help build trust in the agent’s ability to perform reliably in real-world scenarios while maintaining compliance with policy constraints.
@@ 9.3 Evaluation Platforms
    • Dynamic Content: Web applications often involve dynamic elements generated through JavaScript, AJAX, or similar technologies, requiring agents to handle asynchronous updates effectively.
    • Diverse Frameworks: The variety of web technologies (e.g., HTML, CSS, JavaScript frameworks) demands robust agents capable of interacting with a range of interface designs and structures.
@@
TABLE XXIII: Overview of web GUI agent benchmarks (Part I).
@@
Benchmark Platform Year Highlight Data Size Metric Measurement Link
@@
MiniWoB++ [139, 140] Web 2017 Evaluates agents on basic web interactions like clicking, typing, and form navigation. 100 web interaction tasks Task Success Rate Element Match https://github.com/Farama%2DFoundation/miniwob%2Dplusplus
@@
RUSS [136] Web 2021 Uses ThingTalk for mapping natural language to web actions, enabling precise web-based task execution in real HTML environments. 741 instructions Task Success Rate Text Match, Element Match https://github.com/xnancy/russ
@@
WebShop [366] Web 2022 Simulates e-commerce navigation with real-world products, challenging agents with instruction comprehension, multi-page navigation, and strategic exploration. 12,087 instructions Task Success Rate, Step Success Rate" Text Match https://webshop-pnlp.github.io/
@@
Mind2Web [206] Web 2023 Tests adaptability on real-world, dynamic websites across domains. 2,000 tasks Step Success Rate, Task Success Rate Element Match, Action Match https://github.com/OSU-NLP-Group/Mind2Web
@@
Mind2Web-Live [300] Web 2024 Provides intermediate action tracking for realistic task assessment, along with an updated Mind2Web-Live dataset and tools for annotation. 542 tasks Step Success Rate, Task Success Rate, Efficiency Score Element Match, Text Match, trajectory length https://huggingface.co/datasets/iMeanAI/Mind2Web-Live
@@
Mind2Web-Live-Abstracted [256] Web 2024 Abstract the descriptions by omitting task-specific details and user input information in Mind2Web-Live, which are more streamlined and less time-consuming to compose. 104 samples Task Success Rate, Efficiency Score Text Match, Image Match, Element Match, Path Length https://anonymous.4open.science/r/naviqate
@@
WebArena [350] Web 2023 Simulates realistic, multi-tab browsing on Docker-hosted websites, focusing on complex, long-horizon tasks that mirror real online interactions. 812 long-horizon tasks Step Success Rate Text Match https://webarena.dev/
@@
VisualWebArena [351] Web 2024 Assesses multimodal agents on visually grounded tasks, requiring both visual and textual interaction capabilities in web environments. 910 tasks Step Success Rate Text Match, Image Match https://jykoh.com/vwa
@@
MT-Mind2Web [352] Web 2024 Introduces conversational web navigation with multi-turn interactions, supported by a specialized multi-turn web dataset. 720 sessions/3525 instructions Step Success Rate, Turn Success Rate Element Match, Action Match https://github.com/magicgh/self-map
TABLE XXIV: Overview of web GUI agent benchmarks (Part II).
@@
Benchmark Platform Year Highlight Data Size Metric Measurement Link
@@
MMInA [353] Web 2024 Tests multihop, multimodal tasks on real-world websites, requiring agents to handle cross-page information extraction and reasoning for complex tasks. 1,050 tasks Step Success Rate, Task Success Rate Text Match, Element Match https://mmina.cliangyu.com/
@@
AutoWebBench [248] Web 2024 Bilingual web browsing benchmark with 10,000 browsing traces, supporting evaluation across language-specific environments. 10,000 traces Step Success Rate, Efficiency Score Element Match, Action Match, Time https://github.com/THUDM/AutoWebGLM
@@
WorkArena [360] Web 2024 Focuses on real-world enterprise software interactions, targeting tasks frequently performed by knowledge workers 19,912 unique task instances Task Success Rate, Efficiency Score, Completion under Policy, Turn Success Rate Element Match, Text Match, Execution-based Validation https://github.com/ServiceNow/WorkArena
@@
VideoWebArena [361] Web 2024 Focuses on long-context multimodal agents using video tutorials for task completion 74 videos amounting to approximately 4 hours, with 2,021 tasks in total Task Success Rate, Intermediate Intent Success Rate, Efficiency Scores Element Match, State Information, Exact and Fuzzy Text Matches https://github.com/ljang0/videowebarena
@@
EnvDistraction [362] Web 2024 Evaluates the “faithfulness” of multimodal GUI agents by assessing their susceptibility to environmental distractions, such as pop-ups, fake search results, or misleading recommendations 1,198 tasks Task Success Rate Text Match, Element Match, State Information https://github.com/xbmxb/EnvDistraction
@@
WebVLN-v1 [296] Web 2024 Combines navigation and question-answering on shopping sites, integrating visual and textual content for unified web interaction evaluation. 8,990 paths and and 14,825 QA pairs Task Success Rate, Efficiency Score Element Match, Path Length, Trajectory Length https://github.com/WebVLN/WebVLN
@@
WEBLINX [297] Web 2024 Focuses on conversational navigation, requiring agents to follow multi-turn user instructions in realistic, dialogue-based web tasks. 100k interactions Turn Success Rate Element Match, Text Match, Action Match https://mcgill-nlp.github.io/weblinx/
@@
ST-WebAgentBench [354] Web 2024 Evaluates policy-driven safety in web agents, using the Completion under Policy metric to ensure compliance in enterprise-like environments. 235 tasks Task Success Rate, Completion under Policy (CuP), Risk Ratio Element Match, Action Match, Text Match https://sites.google.com/view/st-webagentbench/home
@@
CompWoB [355] Web 2023 Tests agents on sequential, compositional tasks that require state management across multiple steps, simulating real-world automation scenarios. 50 compositional tasks Task Success Rate Element Match https://github.com/google-research/google-research/tree/master/compositional_rl/compwob
TABLE XXV: Overview of web GUI agent benchmarks (Part III).
@@
Benchmark Platform Year Highlight Data Size Metric Measurement Link
@@
TURKING BENCH [356] Web 2024 Uses natural HTML tasks from crowdsourcing to assess interaction skills with real-world web layouts and elements. 32.2K instances Task Success Rate Text Match, Element Match, Image Match https://turkingbench.github.io
@@
VisualWebBench [207] Web 2024 Provides a fine-grained assessment of multimodal large language models (MLLMs) on web-specific tasks 1,534 instances from 139 real websites across 87 sub-domains Task Success Rate, Turn Success Rate, Efficiency Metrics Text Match, Image Match, Element Match, Action Match https://visualwebbench.github.io/
@@
WONDERBREAD [363] Web 2024 Focuses on business process management (BPM) tasks like documentation, knowledge transfer, and process improvement 2,928 human demonstrations across 598 distinct workflows Task Success Rate, Step Success Rate, Efficiency Score, Completion under Policy Text Match, Action Match, State Information https://github.com/HazyResearch/wonderbread
@@
WebOlympus [364] Web 2024 An open platform for web agents that simplifies running demos, evaluations, and data collection for web agents on live websites 50 tasks Task Success Rate, Step Success Rate Action Match /
@@
BrowserGym [367] Web 2024 Provides a unified, extensible, and open-source environment for evaluating web agents with standardized APIs and observations. Benchmarks include MiniWoB(++) with 125 tasks, WebArena with 812 tasks, and WorkArena with up to 341 tasks per level. Task Success Rate, Step Success Rate, Turn Success Rate, Efficiency Metrics. Text-based matching and element match. https://github.com/ServiceNow/BrowserGym

### v11

@@ 8 Models for Optimizing LLM-Brained GUI Agents
TABLE XXV: Overview of foundation models for LLM-brained GUI agents.
@@ 8.1 Foundation Models
Foundation models serve as the core of LLM-powered GUI agents, providing the essential capabilities for understanding and interacting with graphical user interfaces. Recent advancements in both close-source and open-source MLLMs have significantly enhanced the potential of GUI agents, offering improvements in efficiency, scalability, and multimodal reasoning. This subsection explores these foundation models, highlighting their innovations, contributions, and suitability for GUI agent applications. For a quick reference, Table XXV presents an overview of the key models and their characteristics.
@@ 8.2 Large Action Models
TABLE XXVI: An overview of GUI-optimized models on web platforms.
@@ 8.3 LAMs for Web GUI Agents
In the domain of web-based GUI agents, researchers have developed specialized LAMs that enhance interaction and navigation within web environments. These models are tailored to understand the complexities of web GUIs, including dynamic content and diverse interaction patterns. We present an analysis of LAMs tailored for web GUI agents in Table XXVI.
@@
TABLE XXVII: An overview of GUI-optimized models on mobile platforms (Part I).
@@
TABLE XXVIII: An overview of GUI-optimized models on mobile platforms (Part II).
@@
VSC-RL [390] Mobile Android AutoUI [302], Gemini-1.5-Pro / Screenshots GUI actions AitW Addresses sparse-reward, long-horizon tasks for RL by autonomously breaking a complicated goal into subgoals https://ai-agents-2030.github.io/VSC-RL
Digi-Q [392] Mobile Android LLaVA-1.5 [377] 7B GUI screenshots GUI actions, Q-values AitW [352] Introduces a VLM-based Q-function for GUI agent training, enabling reinforcement learning without online interactions. https://github.com/DigiRL-agent/digiq
VEM [393] Mobile Android Qwen2VL [232] 7B GUI screenshots GUI actions, Q-values AitW [352] Unlike traditional RL methods that require environment interactions, VEM enables training purely on offline data with a Value Environment Model. https://github.com/microsoft/GUI-Agent-RL
MP-GUI [394] Mobile Android InternViT-300M and InternLM2.5-7B-chat [521] 8B GUI screenshots Natural language output, element grounding, captioning, and semantic navigation 680K mixed-modality dataset Introduces a tri-perceiver architecture that models textual, graphical, and spatial modalities to enhance GUI reasoning https://github.com/BigTaige/MP-GUI
UI-R1 [395] Mobile Android Qwen2.5-VL-3B 3B GUI screenshots Reasoning text and GUI actions ScreenSpo and AndroidControl Introduces a rule-based reinforcement learning approach using GRPO to enhance reasoning and action prediction in GUI tasks with only 136 examples https://github.com/lll6gg/UI-R1
ViMo [396] Mobile Android Pre-trained Stable Diffusion model [522] / Current GUI image, user action (in natural language), GUI text representation GUI text representation of the next state and reconstructed full GUI image (visual prediction of the next screen) Android Control and AITW First GUI world model that predicts future visual GUI states https://ai-agents-2030.github.io/ViMo/
@@ 8.4 LAMs for Mobile GUI Agents
Figure 27: The PPO training process of VEM [393]. Figure adapted from the original paper.
Mobile platforms present unique challenges for GUI agents, including diverse screen sizes, touch interactions, and resource constraints. Researchers have developed specialized LAMs to address these challenges, enhancing interaction and navigation within mobile environments. We present an overview of LAMs tailored for mobile GUI agents in Table XXVII and XXVIII.
@@
Both Digi-Q [392] and VEM [393] investigate the use of offline RL to enhance the performance of GUI agents without requiring direct interaction with the environment. Digi-Q employs temporal-difference learning to train a Q-function offline and derives policies through a Best-of-N selection strategy based on the predicted Q-values. Similarly, VEM introduces an environment-free RL framework tailored for training LLM-powered GUI agents using PPO. It directly estimates state-action values from offline data by fine-tuning with annotated value data from GPT-4o, thereby enabling policy training without real-time execution in a GUI environment. At inference time, only the policy model is utilized. Figure 27 illustrates the overall architecture of VEM. The study further demonstrates that offline RL with structured credit assignment can achieve performance comparable to interactive RL models. Overall, VEM offers a scalable and layout-agnostic approach for training GUI agents while minimizing interaction costs. Both works underscore the potential of offline RL for GUI agent training.
@@
UI-R1 [395] introduces a RL–based training paradigm aimed at enhancing GUI action prediction for multimodal large language models (MLLMs). The resulting model, UI-R1-3B, fine-tunes Qwen2.5-VL-3B using a novel rule-based reward function that jointly evaluates action type correctness and click coordinate accuracy, while also enabling o1-style [524] chain-of-thought (CoT) reasoning through structured <think> tags. UI-R1 relies on only 136 high-quality samples selected via a three-stage filtering strategy. Despite this limited supervision, UI-R1-3B achieves significant improvements on both in-domain and out-of-domain benchmarks. By leveraging Group Relative Policy Optimization (GRPO) [525], the framework aligns policy optimization with the goals of GUI grounding and task execution. UI-R1 establishes a scalable and data-efficient approach for training GUI agents via RL and paves the way for lightweight yet effective agent design. Its methodology has also been successfully extended to cross-platform agents [404, 405], demonstrating strong generalization capabilities.
In addition to action models, ViMo [396] introduces a novel generative visual world model for GUI agents, aimed at improving App agent decision-making by predicting the next GUI state as an image rather than a textual description. A key innovation of ViMo is the Symbolic Text Representation (STR), which replaces GUI text regions with structured placeholders to facilitate accurate and legible text synthesis. This decoupled design allows the system to handle GUI graphics generation using a fine-tuned diffusion model, and text generation through an LLM, thereby achieving high visual fidelity and semantic precision. ViMo significantly boosts both GUI prediction quality and downstream agent performance, with a reported 29.14% relative improvement in GUI generation metrics and enhanced planning accuracy for long-horizon tasks. As a forward simulator, ViMo represents a crucial advancement toward reliable world models for mobile GUI agents, supporting more effective decision evaluation and trajectory planning in visual environments.
TABLE XXIX: An overview of GUI-optimized models on computer platforms.
@@
ScreenLLM [398] Desktop LLaVA 7B, 13B GUI screenshots Predicted GUI actions High-resolution YouTube tutorials Introduces a novel stateful screen schema to compactly represent GUI interactions over time, enabling fine-grained understanding and accurate action prediction /
@@ 8.5 LAMs for Computer GUI Agents
For desktop and laptop environments, GUI agents must handle complex applications, multitasking, and varied interaction modalities. Specialized LAMs for computer GUI agents enhance capabilities in these settings, enabling more sophisticated task execution. We overview of LAMs for computer GUI agents across in Table XXIX.
@@
TABLE XXX: An overview of GUI-optimized models on cross-platform agents (Part I).
@@
SpiritSight [378] Web, Android, Windows Desktop InternVL [373] 2B, 8B, and 26B GUI screenshots GUI actions AitW [352], CommonCrawl websites, and custom annotations Introduces a Universal Block Parsing (UBP) method to resolve positional ambiguity in high-resolution visual inputs. https://hzhiyuan.github.io/SpiritSight-Agent
TABLE XXXI: An overview of GUI-optimized models on cross-platform agents (Part II).
@@
Magma [403] Web, Mobile, Desktop, Robotics LLaMA-3-8B [90], ConvNeXt-Xxlarge [530] 8.6B GUI screenshots, textual task descriptions GUI actions, robotic manipulation UI, robotics data, human instructional videos Jointly trains on heterogeneous datasets, enabling generalization across digital and physical tasks https://microsoft.github.io/Magma/
GUI-R1 [404] Windows, Linux, MacOS, Android, and Web QwenVL2.5 [531] 3B and 7B GUI screenshots Reasoning text and GUI actions Mixture of 3K high-quality samples first framework to apply rule-based reinforcement learning (RFT) to high-level GUI tasks across platforms. https://github.com/ritzz-ai/GUI-R1.git
InfiGUI-R1 [405] Web, Desktop, and Android Qwen2.5-VL-3B-Instruct 3B GUI screenshots, Accessibility Tree Reasoning text and GUI actions Diverse dataset mixture Two-stage training framework Actor2Reasoner: (1) Reasoning Injection via Spatial Reasoning Distillation, and (2) Deliberation Enhancement via Reinforcement Learning with Sub-goal Guidance and Error Recovery Scenario Construction https://github.com/Reallm-Labs/InfiGUI-R1
Task Generalization [383] Web and Android (Mobile) Qwen2-VL-7B-Instruct [232] 7B GUI screenshots Thoughts and grounded coordinate-based actions 11 domain datasets with 56K GUI trajectory samples Introduces mid-training on diverse non-GUI reasoning tasks (particularly math and code) to substantially enhance GUI agent planning capabilities https://github.com/hkust-nlp/GUIMid
@@ 8.6 Cross-Platform Large Action Models
To achieve versatility across various platforms, cross-platform LAMs have been developed, enabling GUI agents to operate seamlessly in multiple environments such as mobile devices, desktops, and web interfaces. We provide an analysis of LAMs tailored for cross-platform GUI agents in Table XXX and XXXI.
@@
Magma [403] is a foundation model for multimodal AI agents that integrates LLMs with vision and action understanding to complete UI navigation and robotic manipulation tasks. Unlike previous models optimized for either UI automation or robotics, Magma jointly trains on a heterogeneous dataset (about 39M samples) spanning UI screenshots, web navigation, robot trajectories, and instructional videos. It employs SoM and Trace-of-Mark techniques, which enhance action grounding and prediction by labeling actionable elements in GUI environments and tracking motion traces in robotic tasks.
@@ 8.7 Takeaways
  5. 5.
    Inference-Time Computing and Reasoning Models: Recent work highlights the importance of inference-time computing, where models plan, reason, and decompose tasks on the fly without architectural changes. Techniques such as extended context windows and chain-of-thought prompting (e.g., “o1-style” reasoning) enable more robust, long-horizon decision-making. UI-R1 [395], GUI-R1 [404] and InfiGUI-R1 [405] are pioneering efforts in this direction. There is also growing interest in rule-based rewards and cost functions that guide inference-time behavior, integrating explicit heuristics to improve the stability, interpretability, and generalization of GUI agents.
@@ 9 Evaluation for LLM-Brained GUI Agents
In the domain of GUI agents, evaluation is crucial for enhancing both functionality and user experience [56, 58] and should be conducted across multiple aspects. By systematically assessing these agents’ effectiveness across various tasks, evaluation t only gauges their performance in different dimensions but also provides a framework for their continuous improvement [532]. Furthermore, it encourages in vation by identifying areas for potential development, ensuring that GUI agents evolve in tandem with advancements in LLMs and align with user expectations.
@@ 9.1 Evaluation Metrics
Evaluating GUI agents requires robust and multidimensional metrics to assess their performance across various dimensions, including accuracy, efficiency, and compliance (e.g., safety). In a typical benchmarking setup, the GUI agent is provided with a natural language instruction as input and is expected to auto mously execute actions until the task is completed. During this process, various assets can be collected, such as the sequence of actions taken by the agent, step-wise observations (e.g., DOM or HTML structures), screenshots, runtime logs, final states, and execution time. These assets enable evaluators to determine whether the task has been completed successfully and to analyze the agent’s performance. In this section, we summarize the key evaluation metrics commonly used for benchmarking GUI agents. te that different research works may use different names for these metrics, but with similar calculations. We align their names in this section.
@@
    Task Success Rate: Task success rate measures the successful task completion over all tasks set in the benchmark. It evaluates whether the final task completion state is achieved while ig ring the intermediate steps. This metric provides an overall measure of end-to-end task completion, reflecting the agent’s ability to handle complex workflows holistically [419, 431, 413].
@@ 9.2 Evaluation Measurements
    • Partial or Fuzzy Match: Allows for approximate matches, which are useful for handling mi r variations such as typos or sy nyms.
@@
    • System Controls and APIs: Ensuring correct interaction with operating system controls or APIs, such as file dialogs, system menus, or tifications in desktop environments.
@@
Each of these measurement techniques contributes to a comprehensive evaluation framework, ensuring that the agent t only completes tasks but does so with precision, efficiency, and adaptability. Together, they help build trust in the agent’s ability to perform reliably in real-world scenarios while maintaining compliance with policy constraints.
@@ 9.3 Evaluation Platforms
    • Dynamic Content: Web applications often involve dynamic elements generated through JavaScript, AJAX, or similar tech logies, requiring agents to handle asynchro us updates effectively.
    • Diverse Frameworks: The variety of web tech logies (e.g., HTML, CSS, JavaScript frameworks) demands robust agents capable of interacting with a range of interface designs and structures.
@@
TABLE XXXII: Overview of web GUI agent benchmarks (Part I).
@@
Benchmark Platform Year Live Highlight Data Size Metric Measurement Link
@@
MiniWoB++ [144, 145] Web 2017 Yes Evaluates agents on basic web interactions like clicking, typing, and form navigation. 100 web interaction tasks Task Success Rate Element Match https://github.com/Farama%2DFoundation/miniwob%2Dplusplus
@@
RUSS [141] Web 2021 No Uses ThingTalk for mapping natural language to web actions, enabling precise web-based task execution in real HTML environments. 741 instructions Task Success Rate Text Match, Element Match https://github.com/xnancy/russ
@@
WebShop [419] Web 2022 Yes Simulates e-commerce navigation with real-world products, challenging agents with instruction comprehension, multi-page navigation, and strategic exploration. 12,087 instructions Task Success Rate, Step Success Rate" Text Match https://webshop-pnlp.github.io/
@@
Mind2Web [213] Web 2023 No Tests adaptability on real-world, dynamic websites across domains. 2,000 tasks Step Success Rate, Task Success Rate Element Match, Action Match https://github.com/OSU-NLP-Group/Mind2Web
@@
Mind2Web-Live [343] Web 2024 Yes Provides intermediate action tracking for realistic task assessment, along with an updated Mind2Web-Live dataset and tools for an tation. 542 tasks Step Success Rate, Task Success Rate, Efficiency Score Element Match, Text Match, trajectory length https://huggingface.co/datasets/iMeanAI/Mind2Web-Live
@@
Mind2Web-Live-Abstracted [278] Web 2024 Yes Abstract the descriptions by omitting task-specific details and user input information in Mind2Web-Live, which are more streamlined and less time-consuming to compose. 104 samples Task Success Rate, Efficiency Score Text Match, Image Match, Element Match, Path Length https://anonymous.4open.science/r/naviqate
@@
WebArena [406] Web 2023 Yes Simulates realistic, multi-tab browsing on Docker-hosted websites, focusing on complex, long-horizon tasks that mirror real online interactions. 812 long-horizon tasks Step Success Rate Text Match https://webarena.dev/
@@
VisualWebArena [407] Web 2024 Yes Assesses multimodal agents on visually grounded tasks, requiring both visual and textual interaction capabilities in web environments. 910 tasks Step Success Rate Text Match, Image Match https://jykoh.com/vwa
@@
MT-Mind2Web [408] Web 2024 No Introduces conversational web navigation with multi-turn interactions, supported by a specialized multi-turn web dataset. 720 sessions/3525 instructions Step Success Rate, Turn Success Rate Element Match, Action Match https://github.com/magicgh/self-map
TABLE XXXIII: Overview of web GUI agent benchmarks (Part II).
@@
Benchmark Platform Year Live Highlight Data Size Metric Measurement Link
@@
MMInA [409] Web 2024 Yes Tests multihop, multimodal tasks on real-world websites, requiring agents to handle cross-page information extraction and reasoning for complex tasks. 1,050 tasks Step Success Rate, Task Success Rate Text Match, Element Match https://mmina.cliangyu.com/
@@
AutoWebBench [270] Web 2024 No Bilingual web browsing benchmark with 10,000 browsing traces, supporting evaluation across language-specific environments. 10,000 traces Step Success Rate, Efficiency Score Element Match, Action Match, Time https://github.com/THUDM/AutoWebGLM
@@
WorkArena [414] Web 2024 Yes Focuses on real-world enterprise software interactions, targeting tasks frequently performed by knowledge workers 19,912 unique task instances Task Success Rate, Efficiency Score, Completion under Policy, Turn Success Rate Element Match, Text Match, Execution-based Validation https://github.com/ServiceNow/WorkArena
@@
VideoWebArena [415] Web 2024 Yes Focuses on long-context multimodal agents using video tutorials for task completion 74 videos amounting to approximately 4 hours, with 2,021 tasks in total Task Success Rate, Intermediate Intent Success Rate, Efficiency Scores Element Match, State Information, Exact and Fuzzy Text Matches https://github.com/ljang0/videowebarena
@@
EnvDistraction [416] Web 2024 No Evaluates the “faithfulness” of multimodal GUI agents by assessing their susceptibility to environmental distractions, such as pop-ups, fake search results, or misleading recommendations 1,198 tasks Task Success Rate Text Match, Element Match, State Information https://github.com/xbmxb/EnvDistraction
@@
WebVLN-v1 [341] Web 2024 No Combines navigation and question-answering on shopping sites, integrating visual and textual content for unified web interaction evaluation. 8,990 paths and 14,825 QA pairs Task Success Rate, Efficiency Score Element Match, Path Length, Trajectory Length https://github.com/WebVLN/WebVLN
@@
WEBLINX [342] Web 2024 No Focuses on conversational navigation, requiring agents to follow multi-turn user instructions in realistic, dialogue-based web tasks. 100k interactions Turn Success Rate Element Match, Text Match, Action Match https://mcgill-nlp.github.io/weblinx/
@@
ST-WebAgentBench [410] Web 2024 Yes Evaluates policy-driven safety in web agents, using the Completion under Policy metric to ensure compliance in enterprise-like environments. 235 tasks Task Success Rate, Completion under Policy (CuP), Risk Ratio Element Match, Action Match, Text Match https://sites.google.com/view/st-webagentbench/home
@@
CompWoB [411] Web 2023 No Tests agents on sequential, compositional tasks that require state management across multiple steps, simulating real-world automation scenarios. 50 compositional tasks Task Success Rate Element Match https://github.com/google-research/google-research/tree/master/compositional_rl/compwob
TABLE XXXIV: Overview of web GUI agent benchmarks (Part III).
@@
Benchmark Platform Year Live Highlight Data Size Metric Measurement Link
@@
TURKING BENCH [412] Web 2024 Yes Uses natural HTML tasks from crowdsourcing to assess interaction skills with real-world web layouts and elements. 32.2K instances Task Success Rate Text Match, Element Match, Image Match https://turkingbench.github.io
@@
VisualWebBench [214] Web 2024 No Provides a fine-grained assessment of multimodal large language models (MLLMs) on web-specific tasks 1,534 instances from 139 real websites across 87 sub-domains Task Success Rate, Turn Success Rate, Efficiency Metrics Text Match, Image Match, Element Match, Action Match https://visualwebbench.github.io/
@@
WONDERBREAD [417] Web 2024 No Focuses on business process management (BPM) tasks like documentation, knowledge transfer, and process improvement 2,928 human demonstrations across 598 distinct workflows Task Success Rate, Step Success Rate, Efficiency Score, Completion under Policy Text Match, Action Match, State Information https://github.com/HazyResearch/wonderbread
@@
WebOlympus [418] Web 2024 Yes An open platform for web agents that simplifies running demos, evaluations, and data collection for web agents on live websites 50 tasks Task Success Rate, Step Success Rate Action Match /
@@
BrowserGym [420] Web 2024 Yes Provides a unified, extensible, and open-source environment for evaluating web agents with standardized APIs and observations. Benchmarks include MiniWoB(++) with 125 tasks, WebArena with 812 tasks, and WorkArena with up to 341 tasks per level. Task Success Rate, Step Success Rate, Turn Success Rate, Efficiency Metrics. Text-based matching and element match. https://github.com/ServiceNow/BrowserGym